# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ Abstract
The global 6G vision has taken its shape after years of international research and development efforts. This work culminated in ITU-R’s Recommendation on ”IMT-2030 Framework”. While the definition phase of technological requirements is currently ongoing, 3GPP’s standardization process on 6G networks is expected to start in 2025 and worldwide commercialization around 2030. It is timely to present an up-to-date overview of the 6G research field. This article serves as a comprehensive guide to 6G by providing an overall vision, a contemporary survey of the main literature, and an informative tutorial-type presentation style. In our vision, 6G will be based on three fundamental elements: wireless, artificial intelligence (AI), and the Internet of Everything (IoE). Consequently, 6G can ultimately become the Intelligent Network of Everything while serving as an enabling platform for the next major disruption in mobile communication, called mobile intelligence. The potential of mobile intelligence is that anything can be made connected, intelligent, and aware of its environment. This will revolutionize the way how devices, systems, and applications are designed; how they operate and interact with humans and each other; and how they can be used for the benefit of people, society, and the world in general. After high-level visioning, the main details of 6G are discussed, including fundamental elements, disruptive applications, key use cases, main performance requirements, potential technologies, and defining features. A special focus is given to a comprehensive set of potential 6G technologies, each of which is introduced in a tutorial manner, with a discussion on the vision, introduction, past, present, opportunities, challenges, literature, and future research directions. Finally, we speculate on what comes after 6G and sketch the first high-level vision of 7G. All in all, the objective of this article is to provide a thorough guide to 6G in order to serve as a source of knowledge and inspiration for further research and development work in academia, industry, and standardization bodies.
@@ I Introduction
This section takes a glance at the past, present, and future of mobile communications by introducing the evolution from 1G to 6G. After that, a special attention is given on 5G by providing a brief overview on its fundamentals, background, and standardization. The focus is then shifted to 6G. A generic development process and its timeline is discussed first. Then, an up-to-date review is given on the worldwide research activities toward 6G, followed by a contemporary survey of the main 6G vision and survey articles. Finally, the contributions of the article are presented.
@@ I-A A Brief History of Mobile Communications
The generations 1G to 6G are briefly reviewed, with a focus on the disruptive nature of each generation. This evolution is summarized in Figure 1.
@@
The real disruption occurred in the 1990s when 2G networks were introduced. 2G was based on digital communication, allowing calling, texting, and limited data services. Due to digital signal processing technology, transceivers became more complex, powerful, and energy efficient, making phones fit in pockets and having great voice quality. The dominant 2G system was the Global System for Mobile Communications (GSM), which was first used in Europe and later around the world. Other 2G systems were Digital AMPS and cdmaOne in North America and Personal Digital Cellular (PDC) in Japan. In the 2G era, the game changer was the evolution from analog to digital technology, making mobile phones affordable, small, and easy to carry. People were able to call and text anyone anywhere due to high penetration rate and broad coverage. 2G freed the potential of mobile telephony and became a huge success, changing everyday life and interactions between people.
@@
In the 2000s, 3G networks were introduced, with data-centric communication as a key feature. Calling and texting were also included. The radio access technology used in 3G was code-division multiple access (CDMA), allowing for more efficient spectrum use. Turbo coding enabled 3G communication links to approach the Shannon capacity limit in practice. High data rates were possible due to the significantly increased system bandwidth. CDMA2000 was used in North America, while the dominant 3G system elsewhere in the world was the Universal Mobile Telecommunications System (UMTS). 3G provided a platform for the next major disruption, called mobile internet (i.e., mobile broadband). However, mobile internet did not boom in the early phase of 3G since mobile phones did not have touch screens, and their operating systems were not properly optimized for the mobile internet usage. In addition, the majority of Internet content was not optimized for mobile phone usage. In other words, the ecosystem for mobile internet was immature at that time. This led to a clumsy user experience of mobile internet. However, all started to change after the introduction of the iPhone in 2007 and the iPhone 3G in 2008. The iPhones had touch screens and their operating systems were properly optimized for the mobile internet. Touch screens and mobile applications disrupted the whole mobile phone industry. The golden age of the mobile internet, touch screen devices, and mobile applications started at the verge of the 4G era in the late 2000s.
@@
The first commercial 4G networks were launched in 2009. 4G was designed for broadband mobile internet with higher data rates. The air interface technology was changed from 3G’s CDMA to orthogonal frequency-division multiplexing (OFDM). OFDM was chosen due to its robustness, flexibility, and straightforward support for multiple-input multiple-output (MIMO) and lower-complexity receivers. 4G was the first large-scale wireless system to exploit full-dimension MIMO channels. 4G was able to adapt to the prevailing channel and interference conditions. This with the increased system bandwidth and spectrum re-farming enabled high-rate communications and flexible use of broadband mobile internet. The mobile internet boomed in the 4G era and became an essential part of everyday life. The 3rd Generation Partnership Project (3GPP) developed Long-Term Evolution (LTE) (Release 8) was the dominant 4G system commonly used around the world. The evolution of LTE included LTE-Advanced (LTE-A) (Release 10) and LTE-A Pro (Release 13). The main technological components of the evolution were carrier aggregation, enhanced MIMO, relaying, coordinated multi-point (CoMP) transmission/reception, dual connectivity, and licensed-assisted access. LTE evolution also introduced support for vehicle-to-everything (V2X) communications and Machine-Type Communications (MTC).
@@
In 2019, the first 5G networks, based on 3GPP’s 5G System technology (Release 15 [1]), were rolled out. 5G was designed for software-native networking, service-oriented architecture, enhanced mobile internet, mission-critical communications, and massive Internet of Things (IoT) connectivity. In general, 5G is a flexible network with greatly expanded capabilities, providing a robust platform for novel wireless services and applications, particularly for vertical industries. The 5G System introduced a new core network (i.e., 5G Core) and a new radio interface (i.e., 5G New Radio (NR)) as part of the access network. The 5G Core supports service-based architecture (SBA), software-defined networking (SDN), network function virtualization (NFV), and network slicing. The key technologies of 5G radio access network (RAN) include millimeter wave (mmWave) communications, massive MIMO, small cells, flexible OFDM-based transmission scheme, cellular/industrial IoT, edge computing, and private networks. 5G also introduced support for non-terrestrial networks (NTNs), integrated access and backhaul (IAB), NR position, unlicensed NR, unmanned aerial systems (UAS), and NR V2X. Currently, there are four 5G Releases published in 3GPP, i.e., Releases 15 (2018), 16 (2020), 17 (2022), and 18 (2023). The most disruptive feature of the latest Release 18 (i.e., 5G-Advanced) is the introduction of artificial intelligence/machine learning (AI/ML) to mobile networks. This is the turning point when mobile meets intelligence in the network design.
@@
Commercial 6G networks are expected to be launched around 2030. Currently, 6G is in the development phase and standardization will start in 2025. Key technologies under development include terahertz (THz) communications, ultra-massive MIMO, reconfigurable intelligent surfaces (RISs), ultra-dense networks (UDNs), network/edge intelligence, integrated non-terrestrial and terrestrial networks (INTNs), integrated sensing and communication (ISAC), 6G V2X communications, cellular-connected unmanned aerial vehicle (UAV) communications, and THz IAB. Ultimately, 6G aims to be ”the Intelligent Network of Everything”, with extreme performance (capacity, rates, latency, reliability, connection density, coverage, and mobility), pervasive AI/ML (core, edge, and air interface), ubiquitous Internet of Everything (IoE) (sensors, devices, machines, vehicles, drones, robots, etc.), and beyond-communication capabilities (computation, caching, sensing, positioning, and energy). The 6G ecosystem is expected to provide a fruitful platform for the next major disruption, known as mobile intelligence (also known as wireless intelligence and connected intelligence). Mobile intelligence will revolutionize the design, operation, and interactions of devices, systems, and applications. In the big picture, mobile intelligence will be deeply integrated into the future society, providing diverse benefits at all its levels.
@@
This section presents a brief overview of 5G. The main aspects are reviewed, including the introduction, background, and standardization. More comprehensive overviews of 5G have been given in the following publications: 5G NR [2, 3], 5G Evolution [4, 5, 6], and 5G-Advanced [7, 8, 9].
@@ I-B2 Background
The research on 5G began in the early 2010s in academia and industry. At that time, ITU-R also started its development work toward IMT-2020. The role of ITU-R was to build an overall vision, define performance/service requirements, evaluate candidate radio interface technologies (RITs), and select IMT-2020 compliant (5G) systems by the end of 2020. In 2014, the most cited 5G paper, ”What Will 5G Be?”, was published [12]. In this early phase, this study identified the key elements of 5G, providing important guidelines for further research and development work. The first standardization efforts by 3GPP were taken in 2016. At the end of 2018, 3GPP completed the first standard for the 5G System, i.e., Release 15 [1]. Release 15 was the first phase of 5G. The second phase, Release 16 [13], was finalized in mid-2020. Release 17, a continuation of 5G evolution, was frozen in the first quarter of 2022. Release 18, i.e., 5G-Advanced, was completed by the end of 2023. Currently, Release 19 is in preparation and expected to be finalized by mid-2025. The main features of Releases 15-19 are summarized in Table I.
Releases 15 and 16 completed 3GPP’s IMT-2020 submission. ITU-R’s evaluation and selection process culminated in February 2021, when ITU-R published Recommendation ITU-R M.2150-0 ”Detailed specifications of the radio interfaces of IMT-2020” [14]. This document presented three approved IMT-2020 compliant RITs, i.e., 3GPP 5G set of RITs (5G-SRIT), 3GPP 5G-RIT, and 5Gi. The first technology refers to 5G Non-Stand Alone, whereas the second represents 5G Stand Alone. 5G Non-Stand Alone interacts closely with 4G LTE, especially in terms of initial access and spectrum sharing. This Non-Stand Alone version enabled a faster and smoother launch of 5G. 5G Stand Alone works independently without the aid of 4G LTE. The first commercial 5G networks were launched during the first half of 2019. Most of the first 5G deployments were based on the Non-Stand Alone version operating at 3.5 gigahertz (GHz) carrier frequency.
@@ I-B3 Release 15: 5G Phase 1
Release 15 (7/2017-12/2018) introduced the first phase of 5G. The corresponding 5G System consists of three main parts: the core network (5G Core), RAN (5G RAN, including NR), and user equipment (UE) [1]. The 5G Core was designed to support flexible functionalities via its cloud-native service-oriented architecture, enabling novel services and business opportunities. The main features of 5G Core are SDN, SBA, NFV, and network slicing [5, 15]. The 5G RAN was designed to support high-speed mobile internet, mission-critical communication, and massive IoT connectivity. The main features of 5G RAN include mmWave communications (24.25–52.6 GHz), massive MIMO, OFDM with flexible numerology, ultra-lean transmission, dynamic time-division duplex (TDD), forward compatibility, and edge computing [16]. In addition to a new set of features, each release specifies enhancements to a set of existing features. In Release 15, the main enhanced features are MTC, IoT, mission critical, and V2X communications. The Release 15 work items are reviewed in [1].
@@ I-B4 Release 16: 5G Phase 2
Release 16 (1/2019-6/2020) introduced the second phase of 5G. The main new features are private networks, industrial/cellular IoT, NR unlicensed, NR positioning, self-organized network (SON) properties, local area network services, wireless-wireline convergence, high-speed train scenario, and maritime communication services [13]. The main enhanced features include URLLC, V2X, mission-critical services, location services, network slicing, and SBA [13]. Further details on the Release 16 work items can be found in [13].
@@ I-B5 Release 17: 5G Evolution
Release 17 (7/2020-3/2022) continued the work on 5G evolution. The main new features include operation at higher mmWave frequencies (52.7–71 GHz), RAN slicing, IAB, UAS, and NTNs [5]. The existing features that are further enhanced include, among others, massive MIMO, cellular/industrial IoT, URLLC, V2X, multicast/broadcast, positioning, network slicing, edge computing, network automation, non-public networks, and wireless-wireline convergence [5]. The work items of Release 17 are discussed in [17].
@@ I-B6 Release 18: 5G-Advanced
Release 18 (4/2022-12/2023) introduced the first standard of 5G-Advanced, representing a significant evolution step for 5G [7]. A major milestone is the introduction of AI/ML technology, which provides intelligent network features for 5G. Other novel features include mobile IAB, support for UAVs, mobile-terminated small data, and a dual multi-subscriber identity module (SIM) [7]. Enhancements are provided for coverage, mobility, MIMO, multicasting/broadcasting, spectrum sharing, multi-carrier, sidelink, and NTNs [7]. An overview of the Release 18 work items can be found in [18].

### v3

@@ Abstract
The global 6G vision has taken its shape after years of international research and development efforts. This work culminated in ITU-R Recommendation on ”IMT-2030 Framework”. While the definition phase of technological requirements is currently ongoing, 3GPP standardization process on 6G networks is expected to start in 2025 and worldwide commercialization around 2029–2030. It is timely to present an up-to-date overview of the entire 6G research field. This article serves as a comprehensive guide to 6G by providing an overall vision, a contemporary survey of the main literature, and an informative tutorial-type presentation style. In our vision, 6G will be based on three fundamental elements: wireless, artificial intelligence, and Internet of Everything. Consequently, 6G can ultimately become the Intelligent Network of Everything while serving as an enabling platform for the next major disruption in mobile communication, called mobile intelligence. The potential of mobile intelligence is that anything can be made connected, intelligent, and aware of its environment. This will revolutionize the way how devices, systems, and applications are designed; how they operate and interact with humans and each other; and how they can be used for the benefit of people, society, and the world in general. After high-level visioning, the main details of 6G are discussed, including fundamental elements, disruptive applications, key use cases, main performance requirements, potential technologies, and defining features. A special focus is given to a comprehensive set of potential 6G technologies, each of which is introduced in a tutorial manner, with a discussion on the vision, introduction, past, present, opportunities, challenges, literature, and future research directions. Finally, we speculate on what comes after 6G and sketch the first high-level vision of 7G. All in all, the objective of this article is to provide a thorough guide to 6G in order to serve as a source of knowledge and inspiration for further research and development work in academia, industry, and standardization bodies.
@@ I INTRODUCTION
This section takes a glance at the past, present, and future of mobile communications by introducing the evolution from 1G to 6G. After that, a special attention is given on 5G by providing a brief overview on its fundamentals, background, and standardization. The focus is then shifted to 6G. A generic development process and its timeline is discussed first. Then, an up-to-date review is given on the worldwide research activities toward 6G. Finally, the contributions of the article are presented. For the convenience of the readers, the content of this article is introduced in Figure 1.
Figure 1: Content of the article.
@@ I-A A BRIEF HISTORY OF MOBILE COMMUNICATIONS
The generations from 1G to 6G are briefly reviewed, with a focus on the disruptive nature of each generation. This evolution is summarized in Figure 2.
@@
The real disruption occurred in the 1990s, when 2G networks were introduced. 2G was based on digital communication, allowing calling, texting, and limited data services. Due to digital signal processing technology, transceivers became more complex, powerful, and energy efficient, making phones fit in pockets and having a great voice quality. The dominant 2G system was the Global System for Mobile Communications (GSM), which was first used in Europe and later around the world. Other 2G systems were Digital AMPS and cdmaOne in North America and Personal Digital Cellular (PDC) in Japan. In the 2G era, the game changer was the evolution from analog to digital technology, making mobile phones affordable, small, and easy to carry. People were able to call and text anyone anywhere due to the high penetration rate and broad coverage of 2G. 2G freed the potential of mobile telephony and became a huge success, changing everyday life and interactions between people.
@@
In the 2000s, 3G networks were introduced, with data-centric communication as a key feature. Calling and texting were also included. The radio access technology used in 3G was code-division multiple access (CDMA), allowing for more efficient spectrum use. Turbo coding enabled 3G communication links to approach the Shannon capacity limit in practice. High data rates were possible due to the significantly increased system bandwidth. CDMA2000 was used in North America, while the dominant 3G system elsewhere in the world was the Universal Mobile Telecommunications System (UMTS). 3G provided a platform for the next major disruption, called mobile internet (i.e., mobile broadband). However, mobile internet did not boom in the early phase of 3G since mobile phones did not have touch screens, and their operating systems were not properly optimized for mobile internet usage. In addition, the majority of Internet content was not optimized for mobile phone usage. In other words, the ecosystem for mobile internet was immature at that time. This led to a clumsy user experience of mobile internet. However, all started to change after the introduction of the iPhone in 2007 and the iPhone 3G in 2008. The iPhones had touch screens, and their operating systems were properly optimized for mobile internet. Touch screens and mobile applications disrupted the whole mobile phone industry. The golden age of mobile internet, touch screen devices, and mobile applications started at the verge of the 4G era in the late 2000s.
@@
The first commercial 4G networks were launched in 2009. 4G was designed for broadband mobile internet with higher data rates. The air interface technology was changed from CDMA to orthogonal frequency-division multiplexing (OFDM). OFDM was chosen due to its robustness, flexibility, and straightforward support for multiple-input multiple-output (MIMO) and lower-complexity receivers. 4G was the first large-scale wireless system to exploit full-dimension MIMO channels. 4G was able to adapt to the prevailing channel and interference conditions. This with the increased system bandwidth and spectrum re-farming enabled high-rate communications and flexible use of broadband mobile internet. Mobile internet boomed in the 4G era and became an essential part of everyday life. The 3rd Generation Partnership Project (3GPP) developed Long-Term Evolution (LTE) (Release 8) was the dominant 4G system commonly used around the world. The evolution of LTE included LTE-Advanced (LTE-A) (Release 10) and LTE-A Pro (Release 13). The main technological components of the evolution were carrier aggregation, enhanced MIMO, relaying, coordinated multi-point (CoMP) transmission/reception, dual connectivity, and licensed-assisted access. LTE evolution also introduced support for vehicle-to-everything (V2X) communications and Machine-Type Communications (MTC).
@@
In 2019, the first 5G networks, based on the 3GPP 5G System technology (Release 15 [1]), were rolled out. 5G was designed for software-native networking, service-oriented architecture, enhanced mobile internet, mission-critical communications, and massive Internet of Things (IoT) connectivity. In general, 5G is a flexible network with greatly expanded capabilities, providing a robust platform for novel wireless services and applications, particularly for vertical industries. The 5G System introduced a new core network (i.e., 5G Core) and a new radio interface (i.e., 5G New Radio (NR)) as part of the access network. The 5G Core supports service-based architecture (SBA), software-defined networking (SDN), network function virtualization (NFV), and network slicing. The key technologies of 5G radio access network (RAN) include millimeter wave (mmWave) communications, massive MIMO, small cells, flexible OFDM-based transmission scheme, cellular/industrial IoT, edge computing, and private networks. 5G also introduced support for non-terrestrial networks (NTNs), integrated access and backhaul (IAB), NR position, unlicensed NR, unmanned aerial systems (UAS), and NR V2X. Currently, there are four 5G Releases published in 3GPP, i.e., Releases 15 (2018), 16 (2020), 17 (2022), and 18 (2023). The most disruptive feature of Release 18 (i.e., 5G-Advanced) was the introduction of artificial intelligence/machine learning (AI/ML) to mobile networks. This was the turning point when mobile met intelligence in the network design.
@@
Commercial 6G networks are expected to be launched around 2029–2030. Currently, 6G is in the development phase and standardization will start in 2025. Key technologies under development include terahertz (THz) communications, ultra-massive MIMO, reconfigurable intelligent surfaces (RISs), ultra-dense networks (UDNs), network/edge intelligence, integrated non-terrestrial and terrestrial networks (INTNs), integrated sensing and communication (ISAC), 6G V2X communications, cellular-connected unmanned aerial vehicle (UAV) communications, and THz IAB. Ultimately, 6G aims to be ”the Intelligent Network of Everything”, with extreme performance (capacity, rates, latency, reliability, connection density, coverage, and mobility), pervasive AI/ML (core, edge, and air interface), ubiquitous Internet of Everything (IoE) (sensors, devices, machines, vehicles, drones, robots, etc.), and beyond-communication capabilities (computation, caching, sensing, positioning, and energy). The 6G ecosystem is expected to provide a fruitful platform for the next major disruption, known as mobile intelligence (also known as wireless intelligence and connected intelligence). Mobile intelligence will revolutionize the design, operation, and interactions of devices, systems, and applications. In the big picture, mobile intelligence will be deeply integrated into the future society, providing diverse benefits at all its levels.
@@
This section presents a brief overview of 5G. The main aspects are reviewed, including introduction, background, and standardization. More comprehensive overviews of 5G have been presented in the following publications: 5G NR [2], 5G Evolution [3, 4, 5], and 5G-Advanced [6, 7, 8].
@@ I-B2 BACKGROUND
The research on 5G began in the early 2010s in academia and industry. At that time, ITU-R also started its development work toward IMT-2020. The role of ITU-R was to build an overall vision, define performance/service requirements, evaluate candidate radio interface technologies (RITs), and select IMT-2020 compliant (5G) systems by the end of 2020. In 2014, a landmark 5G paper, ”What Will 5G Be?”, was published [11]. In this early phase, this study identified the key elements of 5G, providing important guidelines for further research and development work. The first standardization efforts by 3GPP were taken in 2016. At the end of 2018, 3GPP completed the first standard for the 5G System, i.e., Release 15 [1]. Release 15 was the first phase of 5G. The second phase, Release 16 [12], was completed in mid-2020. Release 17, a continuation of 5G evolution, was finalized in the first quarter of 2022. Release 18, i.e., 5G-Advanced, was completed by the end of 2023. Currently, Release 19 is in preparation and expected to be finalized by mid-2025. Release 20 will be the last 5G dominant Release, along with the first studies on 6G. The estimated timeline of Release 20 is 7/2025–12/2026. The main features of Releases 15–19 are summarized in Table I.
Releases 15 and 16 were submitted to the ITU-R evaluation process as a potential RIT candidate to meet the IMT-2020 requirements. The evaluation and selection process culminated in February 2021, when ITU-R published Recommendation ITU-R M.2150-0 ”Detailed specifications of the radio interfaces of IMT-2020” [13]. This document presented three approved IMT-2020 compliant systems, i.e., 3GPP 5G set of RITs (SRITs), 3GPP 5G RIT, and 5G RIT for India (5Gi). The first technology refers to 5G Non-Stand Alone (NSA), whereas the second represents 5G Stand Alone (SA). 5G NSA is dependent on 4G LTE and cannot work independently. The NSA version enabled a faster and smoother launch of 5G due to its easier implementation compared to SA. Unlike NSA, 5G SA operates independently without the aid of 4G LTE. The first commercial 5G networks were launched during the first half of 2019. Most of the first 5G deployments were based on the NSA version, operating at 3.5 gigahertz (GHz) carrier frequency.
@@ I-B3 RELEASE 15: 5G PHASE 1
Release 15 (7/2017–12/2018) introduced the first phase of 5G. The corresponding 5G System comprised three main parts: 5G Core, 5G RAN, and 5G user equipment (UE) [1]. The 5G Core was designed to support flexible functionalities via its cloud-native service-oriented architecture, enabling novel services and business opportunities. The main features of 5G Core were SDN, SBA, NFV, and network slicing [4]. The 5G RAN was designed to support high-speed mobile internet, mission-critical communication, and massive IoT connectivity. The main features of 5G RAN included mmWave communications (24.25–52.6 GHz), massive MIMO, OFDM with flexible numerology, ultra-lean transmission, dynamic time-division duplex (TDD), forward compatibility, and edge computing [14]. In addition to a new set of features, each release specifies enhancements to a set of existing features. In Release 15, the main enhanced features were MTC and V2X communications. A comprehensive review of the Release 15 work items can be found in [1].
@@ I-B4 RELEASE 16: 5G PHASE 2
Release 16 (1/2019–6/2020) introduced the second phase of 5G. The main new features were private networks, industrial/cellular IoT, NR unlicensed, NR positioning, self-organized network (SON) properties, local area network services, wireless-wireline convergence, high-speed train scenario, and maritime communication services [12]. The main enhanced features were URLLC, V2X, mission-critical services, location services, network slicing, and SBA [12]. Further details on the Release 16 work items can be found in [12].
@@ I-B5 RELEASE 17: 5G EVOLUTION
Release 17 (7/2020–3/2022) continued the work on 5G evolution. The main new features were operation at higher mmWave frequencies (52.7–71 GHz), RAN slicing, NR light, IAB, UAS, and NTNs [4]. The existing features that were further enhanced included massive MIMO, cellular/industrial IoT, URLLC, V2X, multicast/broadcast, positioning, network slicing, edge computing, network automation, non-public networks, and wireless-wireline convergence [4]. More information on the work items of Release 17 can be found in [15].
@@ I-B6 RELEASE 18: 5G-ADVANCED
Release 18 (4/2022–12/2023) introduced the first standard of 5G-Advanced, representing a significant evolution step for 5G [6]. A major milestone was the introduction of AI/ML technology, which provides intelligent network features for 5G. Other novel features were mobile IAB, mobile-terminated small data, and a dual multi-subscriber identity module (SIM) [6]. Enhancements were provided for coverage, mobility, MIMO, multicasting/broadcasting, spectrum sharing, multi-carrier, sidelink, and NTNs [6]. An overview of the Release 18 work items can be found in [16].